ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 264, 2024

Basu, R. and M. Bhatia Small deviation estimates and small ball probabilities for geodesics in last passage percolation

37

Becher, K. J., N. Grenier-Boley and J.-P. Tignol The discriminant Pfister form of an algebra with involution of capacity four

235

Bhatia, M. See R. Basu Cohen, G. and M. Lin L^{2}-Quasi-compact and hyperbounded

Markov operators

281

De Stefani, A. and I. Smirnov Stability and deformation of F - singularities

1

Fernandez Herrero, A. On the quasicompactness of the moduli stack of logarithmic G-connections over a curve

389

Grenier-Boley, N. See K. J. Becher Kamčev, N., A. Liebenau and N. Morrison On uncommon sys-

tems of equations

331

Krepski, D., J. Watts and S. Wolbert Sheaves, principal bundles, and Čech cohomology for diffeological spaces

461

Liebenau, A. See N. Kamčev Lin, M. See G. Cohen Maglione, J. and C. Voll Flag Hilbert–Poincaré series and Igusa

zeta functions of hyperplane arrangements

177

Monden, N. On minimal generating sets for the mapping class group of a punctured surface

97

Morrison, N. See N. Kamčev Mountakis, A. Distinguishing sets of strong recurrence from van der

Corput sets

121

Shestakov, I. and S. Sverchkov New central elements in free alternative algebras

363

Smirnov, I. See A. De Stefani Sverchkov, S. See I. Shestakov Tignol, J.-P See K. J. Becher Tsodikovich, D. An analogue of the Blaschke–Santaló inequality for

billiard dynamics

429

Voll, C. See J. Maglione Watts, J. See D. Krepski Wolbert, S. See D. Krepski Zhang, H. and R. Zhang Nondivergence on homogeneous spaces

and rigid totally geodesic submanifolds

149

Zhang, R. See H. Zhang